# DualFluidNet: an Attention-based Dual-pipeline Network for Accurate and Generalizable Fluid-solid Coupled Simulation

### v1

@@ Abstract
Fluid motion can be considered as point cloud transformation when adopted by a Lagrangian description. Compared to traditional numerical analysis methods, using machine learning techniques to learn physics simulations can achieve near accuracy, while significantly increasing efficiency. In this paper, we propose an innovative approach for 3D fluid simulations utilizing an Attention-based Dual-pipeline Network, which employs a dual-pipeline architecture, seamlessly integrated with an Attention-based Feature Fusion Module. Unlike previous single-pipeline approaches, we find that a well-designed dual-pipeline approach achieves a better balance between global fluid control and physical law constraints. Furthermore, we design a Type-aware Input Module to adaptively recognize particles of different types and perform feature fusion afterward, such that fluid-solid coupling issues can be better dealt with. The experiments show that our approach significantly increases the accuracy of fluid simulation predictions and enhances generalizability to previously unseen scenarios. We demonstrate its superior performance over the state-of-the-art approaches across various metrics.
@@ 1 Introduction
Comprehending physics is essential for enhancing our understanding of the environment and our interactions with it. Traditional fluid simulations have limitations in practical applications due to their high computational demands. Recent studies [29, 3, 22, 17, 26] have highlighted the potential of physics-based neural networks as a promising approach for comprehending intricate natural phenomena from data. Inspired by Smoothed Particle Hydrodynamics (SPH) methods [6, 11], which represent fluids as smoothed particles, several networks [13, 19, 24, 27, 8] have been developed for particle-based simulations. The prediction of physical properties using neural networks is always treated as a regression problem, enabling the effective learning and approximation of physical phenomena.
@@
While previous methods have demonstrated excellent performance in fluid simulation, we have observed that single-path networks always struggle to maintain a balance between overall particle control accuracy and adherence to physical laws. The dilemma in single-path methods turns into a trade-off issue. To deal with this problem, we propose a Dual-Pipeline Network architecture, consisting of a main pathway and a constraint-guided pathway. The main pathway is responsible for establishing a global context and stabilizing the fluid. The constraint-guided pathway incorporates enforced constraints, learning fluid motion patterns under the physical laws. Our network seamlessly fuses the two learning ways through a meticulously designed approach, achieving a harmonious balance between these two pathways and unlocking their maximum potential. Instead of manually determining the proportion of importance for the two pathways, we employ an innovative Attention-based Feature Fusion Module, allowing the network to autonomously learn the optimal fusion between the two paths.
@@
Although previous networks appear to perform well in handling fluid point clouds, there hasn’t been an effective method for handling the mixed input of fluid and solid. Common practices involve simply concatenating the two inputs and feeding them into the network [27], or passing each through a few convolutional layers before concatenation [19]. These straightforward input concatenations are insufficient to enable the network to effectively distinguish between the features of fluid and solid particles. Nevertheless, the fluid-solid coupling is crucial for enhancing fluid simulation performance. Common feature fusion methods are frequently utilized for multimodal and multiscale fusion problems [31, 15, 5]. In the context of fluid simulation, our input comprises two entirely different features: liquid particles and solid particles. Additionally, the particle count varies in each input. We propose an innovative approach to facilitate better differentiation between fluid and solid particles in the input, enabling the network to more accurately and reliably compute their interactions during collisions. We verified that the Type-aware Input Module can be easily transferred to other network models and plays a crucial role in enhancing fluid simulation performance.
@@
We conducted experiments on multiple 3D liquid datasets, and the results demonstrated that our network outperformed existing methods across various performance metrics, surpassing even the state-of-the-art network [19]. Furthermore, we validated the strong generalization of our network model in more challenging scenarios.
@@ 2 Related work
Comprehending physics is crucial for making sense of our surroundings and interacting with them more effectively. Smoothed Particle Hydrodynamic (SPH), representing fluids as smoothed particles, is a meshless Lagrangian method commonly employed for simulating fluid dynamics. This perspective inherently conserves fluid mass since the particle count remains constant and efficiently focuses computational resources on the fluid’s location. Traditional fluid simulation methods exhibit a high level of simulation fidelity [25, 2, 30, 32, 12], but they often entail complex computations and significant computational expenses, thereby constraining their utility in situations that require real-time predictions. In contrast, neural networks have emerged as a highly promising approach to capturing the intricacies of natural phenomena from data without incurring enormous computational costs [14, 26, 17].
A crucial consideration in employing neural networks for physics-based learning is selecting the appropriate data representation. Point clouds are widely employed in computer vision and robotics applications to represent the 3D environment[20, 21, 28, 4]. PointNet[20], which directly applies deep learning models to point cloud data, is an excellent feature extractor. However, for complex and dynamic 3D fluid simulations, a more sophisticated approach is required for accurate computation. A common approach is to utilize graph structures [13, 1, 24, 18] to create a dynamic graph connecting nearby particles, facilitating the modeling of fluid behavior. This is achieved by updating nodes and edges to simulate ’message passing’ between SPH particles. Graph-based methods are strongly tied to particle discretization, but a wide range of physical phenomena, such as fluid mechanics, are governed by continuous partial differential equations (PDE) instead of discrete graph structures.
@@
Most related to our approach is [27, 19]. [27] proposed a continuous convolution kernel (CConv) designed for establishing connections between particles and their neighbors. CConv exhibits strong learning capabilities for fluid particles and can broadly capture the dynamics of fluid motion. However, due to the absence of explicit physical constraints, its predictions may deviate from physical laws. Therefore, [19] has ingeniously designed an anti-symmetric convolution kernel building upon the continuous convolution kernel from CConv to introduce a robust constraint, allowing the network to approximate the conservation of momentum in physics during training. However, this strong constraint significantly complicates the problem for the network, leading to a reduction in its learning capacity. Single-path networks always confront this dilemma, while our dual-pipeline network excels at balancing these two distinct learning approaches, enabling our network to achieve a better equilibrium between global fluid control and adherence to physical law constraints.
@@ 3 Method
  $\bm{Input:}x^{t}=\left\{x_{1},x_{2},...x_{N}\right\},v^{t}=\left\{v_{1},v_{2},...v_{N}\right\}$
@@
     apply external forces $v_{i}^{*}\Leftarrow v_{i}+\bigtriangleup t\frac{F_{ext}}{m}$
     update position $x_{i}^{*}\Leftarrow x_{i}+\Delta tv_{i}^{*}$
@@
  $x^{t*}=\left\{x_{1}^{*},x_{2}^{*},...x_{N}^{*}\right\}$
  $v^{t*}=\left\{v_{1}^{*},v_{2}^{*},...v_{N}^{*}\right\}$
  predict position $\Delta x^{t+1}=M_{\theta}\left(x^{t*},v^{t*}\right)$
@@
Our network follows the position-based fluids (PBF) [16] scheme. Algorithm 1 outlines the general procedure of this approach, where $M_{\theta}$ represents our neural network model with trainable parameters $\theta$, and $\Delta t$ represents the time interval between two consecutive time steps. Given a set of particles, denoted as $p_{N}$, where $N$ represents the particle number, the algorithm proceeds to analyze their positions and velocities. It takes $x^{t}$ and $v^{t}$ as inputs, where $x^{t}$ represents the positions of the particles at time step $t$ and $v^{t}$ represents their velocities. Our central learning goal is to approximate the underlying physical dynamics to predict the state of particles at the next time step $t+1$.
@@ 3.1 Two Types of Convolution Kernels
We introduce two types of convolution kernels, denoted as CConv[27] and ASCC[19]. Given a point cloud $p_{N}$ containing $N$ points labeled as $i=1,...,N$, with corresponding values $f_{i}$ at positions $x_{i}$, the convolution of CConv at position $x\in p_{n}$ can be defined as follows:
@@
Where $\mathcal{N}\left(x,R\right)$ represents the set of points located within a radius R around the position x. $g$ is the filter function, utilizing a mapping $\Lambda(r)$ that transforms a unit ball into a unit cube, described by [7, 27]. $a$ is a window function applied for density normalization specific to the points $x_{i}$ and $x$ as in [10]:
$a\left(x_{i},x\right)=\left\{\begin{array}[]{lc}\left(1-\frac{\left\|x_{i}-x_{2}^{2}\right\|}{R^{2}}\right)^{3}&\left\|x_{i}-x\right\|_{2}<R\\ 0&else.\\ \end{array}\right.$ (2)
Figure 3: The anti-symmetric kernel is obtained by negating and mirroring the trainable variables by the center point.
@@
Where $g_{s}$ represents the anti-symmetric convolution kernel and $f$ represents the corresponding value at the position $x$, with the remaining symbols defined similarly to the definitions in Equation 1.
3.2 Dual-pipeline Architectrue
Our network architecture is illustrated in Figure 2. Instead of manually determining the proportion of importance for the two pathways, we employ an innovative Attention-based Feature Fusion Module, allowing the network to autonomously learn the optimal connection between the two pathways. Instead of manually assigning specific weights to each pathway, we utilize an innovative Attention-based Feature Fusion Module, shown as Figure 3(a), enabling the network to autonomously learn the optimal fusion between the two pathways.
More precisely, our network comprises a main pathway and a constraint-guided pathway. Both pathways share a similar structure, each comprising a Type-aware Input Module and a sequence of continuous convolutions with an effective depth of five. The main pathway, which utilizes CConv as its kernel, is responsible for establishing a global context and stabilizing the fluid. The Constraint-guided pathway employs Anti-symmetric Convolution as its kernel, which enforces a strong constraint to ensure momentum conservation while learning fluid motion patterns. Before feeding the input into the convolutional sequence, we combine the result from the input module and a fully-connected layers by addition. In the second and fourth layers of the convolution sequence, we leverage the Feature Fusion Module to implement a residual connection concept [9]. We found that these designs improve accuracy and robustness.
@@ 3.2 Dual-pipeline Architectrue
Figure 4: The schematic of the Attention-based Feature Fusion Module. Features from both pathways pass through two layers of CConv and are then concatenated. The concatenated features are subsequently input into a Particle Selector to determine the respective weights for fusion. The original features from both pathways are then combined with weighted fusion.

### v2

# DualFluidNet: an Attention-based Dual-pipeline Network for Fluid Simulation

@@ Abstract
Fluid motion can be considered as a point cloud transformation when using the SPH method. Compared to traditional numerical analysis methods, using machine learning techniques to learn physics simulations can achieve near-accurate results, while significantly increasing efficiency. In this paper, we propose an innovative approach for 3D fluid simulations utilizing an Attention-based Dual-pipeline Network, which employs a dual-pipeline architecture, seamlessly integrated with an Attention-based Feature Fusion Module. Unlike previous methods, which often make difficult trade-offs between global fluid control and physical law constraints, we find a way to achieve a better balance between these two crucial aspects with a well-designed dual-pipeline approach. Additionally, we design a Type-aware Input Module to adaptively recognize particles of different types and perform feature fusion afterward, such that fluid-solid coupling issues can be better dealt with. Furthermore, we propose a new dataset, Tank3D, to further explore the network’s ability to handle more complicated scenes. The experiments demonstrate that our approach not only attains a quantitative enhancement in various metrics, surpassing the state-of-the-art methods, but also signifies a qualitative leap in neural network-based simulation by faithfully adhering to the physical laws. Code and video demonstrations are available at https://github.com/chenyu-xjtu/DualFluidNet.
keywords:
Fluid Simulation, Learning Physics, Neural Network, Deep Learning
journal: Neural Networks
[…]
organization=School of Software Engineering, Xi’an Jiaotong University,city=Xi’an, postcode=710049, country=China
@@ 1 Introduction
(a) Real-time rendered fluid surface
(b) Underlying simulation particles
@@
Comprehending physics is essential for enhancing our understanding of the environment and our interactions with it. Traditional fluid simulations have limitations in practical applications due to their high computational demands. Recent studies [16, 45, 3, 18, 8, 6] have highlighted the potential of physics-based neural networks as a promising approach for comprehending intricate natural phenomena from data. Inspired by Smoothed Particle Hydrodynamics (SPH) methods [9, 19], which represent fluids as smoothed particles, several networks [17, 23, 29, 34, 39, 11] have been developed for particle-based simulations. The prediction of physical properties using neural networks is always treated as a regression problem[20], enabling the effective learning of physical phenomena.
@@
While previous methods have demonstrated excellent performance in fluid simulation by using neural network methods, we have observed that previous methods always struggle to maintain a balance between overall particle control accuracy and adherence to physical laws. The dilemma in previous methods turns into a trade-off issue. This limitation can result in deformations and disintegrations that defy the laws of physics, sometimes even in the simplest cases of free-fall motion. To deal with this problem, we propose a Dual-Pipeline Network architecture, consisting of a main pathway and a constraint-guided pathway. On one hand, the main pathway exhibits stronger learning capabilities of overall fluid dynamics, thereby being responsible for establishing a global context and stabilizing the fluid. On the other hand, although the constraint-guided pathway has relatively weaker overall learning capabilities, it can comprehend fluid particle motion while adhering to physical laws, thereby integrating physical laws into the main pathway. Our network seamlessly fuses the two learning ways through a meticulously designed approach, achieving a harmonious balance between these two pathways and unlocking their maximum potential. Instead of manually deciding how much each of the two pathways contributes, we propose a new module designed specifically for fluid particles, enabling each layer of the network to autonomously determine the optimal feature fusion parameters. This allows the network to automatically learn the best way to combine features from both pathways.
@@
The integration of both fluid and solid inputs is often neglected in neural network-based methods for fluid simulation. Although previous networks appear to perform well in handling fluid point clouds, there hasn’t been an effective method for handling the mixed input of fluid and solid. Common practices involve simply concatenating the two inputs and feeding them into the network [39], or passing each through a few convolutional layers before concatenation [29]. These straightforward input concatenations are insufficient to enable the network to effectively distinguish between the features of fluid and solid particles. Nevertheless, the fluid-solid coupling is crucial for enhancing fluid simulation performance. In the context of fluid simulation, our input comprises two entirely different features: fluid particles and solid particles. Additionally, the particle count varies in each input. Based on this observation, we propose an innovative Type-aware Input Module to facilitate better differentiation between fluid and solid particles in the input, enabling the network to more accurately and reliably compute their interactions during collisions. We verified that this module can be easily transferred to other network models and plays a crucial role in enhancing fluid simulation performance.
@@
We conducted experiments on multiple 3D liquid datasets, and the results demonstrated that our network outperformed existing methods across various performance metrics, surpassing even the state-of-the-art network [29]. We also propose a new 3D liquid dataset, Tank3D, to further explore the network’s ability to handle more complex scenes. Furthermore, we validated the strong generalization of our network model in more challenging scenarios.
@@ 2 Related work
2.1 Representation of Fluid
In Computational Fluid Dynamics (CFD), fluids are typically represented using either the grid-based or particle-based methods, the latter also known as Smoothed Particle Hydrodynamics (SPH). The grid-based method divides the fluid domain into a grid of cells or voxels. Properties such as velocity, pressure, and density are assigned to each grid point. The SPH method represents fluids as discrete particles. Each particle carries properties such as velocity, mass, and density. Properties at specific locations are computed by averaging the attributes of neighboring particles. This process is performed using a kernel function, which assigns weights to nearby particles based on their distances from the target position. The SPH kernel functions produce non-zero values within a fixed radius range and are zero outside this range. Because grid-based methods often struggle with handling complex geometries and regions undergoing substantial deformation, and SPH can better utilize GPU parallel computing capabilities, the SPH approach has been widely adopted for fluid representation in recent years.
@@ 2.1 Representation of Fluid
Traditional SPH methods exhibit a high level of simulation fidelity [36, 2, 46, 49, 21], but they often entail complex computations and significant computational expenses, thereby constraining their utility in situations that require real-time predictions. In contrast, using machine learning to simulate the SPH method has emerged as a highly promising approach to capturing the intricacies of natural phenomena from data without incurring enormous computational costs [24, 38, 27, 33]. Due to the strong demand for generalization and feature extraction capabilities in fluid simulation, deep learning methods have been increasingly favored in recent years over traditional machine learning approaches [20]. In utilizing neural networks for physics-based learning, a crucial consideration is selecting the appropriate data representation. Point clouds are widely employed in computer vision and robotics applications to represent the 3D environment [30, 31, 44, 4]. In the SPH method using neural networks, fluid particles can also be regarded as point clouds, while fluid simulation can be visualized as point cloud transformation.
2.2 Neural Network Approaches in Fluid Simulation
In time series analysis and prediction tasks, RNN-based models such as Long Short-Term Memory (LSTM) or Gated Recurrent Unit (GRU) are commonly employed [22, 7]. This is attributed to the temporal dependencies and sequential nature inherent in time series data, which are effectively captured by recurrent neural network models. While using RNN-based networks to predict the state of fluid particles frame by frame might seem like an intuitive choice, we observe that employing CNN-based networks is more appropriate in practice. By using convolution kernel to simulate the kernel function in SPH method, the state (position and velocity) of fluid particles at timestep $t+1$ can be directly computed from the particle state at timestep $t$, without requiring information from timesteps prior to $t$. Conversely, incorporating the hidden states of RNN-based networks will introduce instability into the computation.
@@ 2.2 Neural Network Approaches in Fluid Simulation
A common convolution approach is to utilize graph structures [35, 23, 1, 34, 28] to create a dynamic graph connecting nearby particles, facilitating the modeling of fluid behavior. This is achieved by updating nodes and edges to simulate message passing between SPH particles. Graph-based methods are tied to particle discretization, but a wide range of physical phenomena, such as fluid mechanics, are governed by continuous partial differential equations (PDE) instead of discrete graph structures. Therefore, continuous methods [43, 39, 29, 37] always exhibit a stronger capability. Most related to our approach is [39, 29]. [39] proposed a continuous convolution kernel (CConv) designed for establishing connections between particles and their neighbors. CConv exhibits strong learning capabilities for fluid particles and can broadly capture the dynamics of fluid motion. However, due to the absence of explicit physical constraints, its predictions may deviate from physical laws. Therefore, [29] has designed an anti-symmetric continuous convolution kernel (ASCC) building upon the CConv to introduce a robust constraint, allowing the network to approximate the conservation of momentum in physics during training. However, this strong constraint also complicates the problem for the network, leading to a reduction in its learning capacity. These networks always face such a dilemma, while our dual-pipeline network effectively integrates the strengths of both distinct learning approaches, enabling our network to achieve a better equilibrium between global fluid control and adherence to physical law constraints.
2.3 Feature Fusion
Common feature fusion methods often appear in problems related to multimodal fusion [48, 25], multi-view image fusion [13, 41], multiscale feature fusion [42, 12, 5, 47], and so on. Different from the scenarios mentioned above, feature fusion in this paper is addressed in two scenarios: the fusion of features from two pathways (Attention-based Feature Fusion), and the blending of fluid particles and solid particles inputs (Type-aware Input Module). As for the Attention-based Feature Fusion, the emphasis is on leveraging the strong learning capability of the main path to capture global fluid features, while using the constraint-guided path to incorporate physical information into the main path. Compared to a single pathway, the fusion of two pathways complements each other and maximizes the advantages of both. And the Type-aware Input Module emphasizes a better understanding and differentiation between two distinct types of input: fluid and solid particles in the input. Its purpose is to enable the network to accurately and reliably compute their interactions during collisions.
Both the Attention-based Feature Fusion and Type-aware Input Module utilize a similar soft-attention mechanism, specifically designed for fluid particles. We also consider other attention mechanisms, particularly self-attention and cross-attention [40]. Both mechanisms operate by computing the dot product between queries (Q), keys (K), and values (V), which are transformation representations of elements in the sequence. Self-attention is especially effective for capturing long-range dependencies within a sequence, while cross-attention helps the model capture alignment relationships between two sequences. These attention mechanisms are also commonly employed in feature fusion tasks, as they enable the model to directly compute interactions between any two positions in the sequence. However, in SPH fluid simulation, the local support property of SPH kernel functions ensures that they only produce non-zero values within a fixed radius range, being zero outside of this range. This implies that each particle is influenced only by nearby particles within this radius, with far particles having little effect. Consequently, SPH simulations do not require self-attention and cross-attention mechanisms to explicitly capture long-range dependencies. Instead, through our designed CConv-based soft-attention feature fusion for fluid simulation, we can better fit the local properties of the kernel function, effectively simulating particle interactions while ensuring higher efficiency.
@@ 3 Method
Before introducing our method, the mathematical notation and preliminary definitions are provided in Table 1.
$F_{ext}$ external force
$N_{\theta}$ neural network with parameters $\theta$
$P_{n}$ a set of $n$ particles
$x_{i}^{t},v_{i}^{t}$ state of particle $i$ at time $t$
$x^{t},v^{t}$ state of particles at time $t$
$x_{i}^{t*},v_{i}^{t*}$ state of particle $i$ at time $t$ after $F_{ext}$
$x^{t*},v^{t*}$ state of particles at time $t$ after $F_{ext}$
$\Delta x^{t+1}$ displacement of particles by internal forces
$x^{t+1},v^{t+1}$ state of predicted particles at time $t+1$
$\hat{x}_{i}^{t+1},\hat{v}_{i}^{t+1}$ state of ground-truth particles at time $t+1$
* convolution operator
$\Lambda$ a mapping function that transforms a unit ball into a unit cube
$\left\|\ \cdot\ \right\|_{2}$ $L2$ norm (or called Euclidean norm)
$\phi$ function of CConv
$\psi$ function of ASCC
$\oplus$ concatenate
$\otimes$ weighted average based on weight $w$
$S$ function of Particle Selector
Table 1: A summary table of the main notations. The ”state of particle” includes the position and velocity.
@@
  $\bm{Input:}x^{t}=\left\{x^{t}_{1},x^{t}_{2},...x^{t}_{n}\right\},v^{t}=\left\{v^{t}_{1},v^{t}_{2},...v^{t}_{n}\right\}$
@@
     apply external forces $v_{i}^{t*}=v^{t}_{i}+\Delta t\frac{F_{ext}}{m}$
     update position $x_{i}^{t*}=x_{i}^{t}+\Delta tv_{i}^{t*}$
@@
  $x^{t*}=\left\{x_{1}^{t*},x_{2}^{t*},...x_{n}^{t*}\right\}$
  $v^{t*}=\left\{v_{1}^{t*},v_{2}^{t*},...v_{n}^{t*}\right\}$
  predict position $\Delta x^{t+1}=N_{\theta}\left(x^{t*},v^{t*}\right)$
@@
Our network follows the position-based fluids (PBF) [26] scheme. Algorithm 1 outlines the general procedure of this approach, where $N_{\theta}$ represents our neural network model with trainable parameters $\theta$, and $\Delta t$ represents the time interval between two consecutive time steps. Given a set of particles, denoted as $P_{n}$, where $n$ represents the particle number, the algorithm proceeds to analyze their positions and velocities. It takes $x^{t}$ and $v^{t}$ as inputs, which represent the positions and velocities of the particles at time step $t$. Our central learning goal is to approximate the underlying physical dynamics to predict the positions and velocities of particles at the next time step $t+1$.
@@ 3.1 Two Types of Convolution Kernels
We introduce two types of convolution kernels, denoted as CConv and ASCC. Given a point cloud $P_{n}$ containing $n$ points labeled as $i=1,...,n$, with corresponding values $f_{i}$ at positions $x_{i}$, the convolution of CConv at position $x\in P_{n}$ can be defined as follows:
@@
Where $\mathcal{N}\left(x,R\right)$ represents the set of points located within a radius R around the position x. $g$ is the filter function, utilizing a mapping $\Lambda(r)$ that transforms a unit ball into a unit cube, described by [10, 39]. $a$ is a window function applied for density normalization [15] specific to the points $x_{i}$ and $x$:
$a\left(x_{i},x\right)=\left\{\begin{array}[]{lc}\left(1-\frac{\left\|x_{i}-x\right\|_{2}^{2}}{R^{2}}\right)^{3}&\left\|x_{i}-x\right\|_{2}<R\\ 0&else.\\ \end{array}\right.$ (2)
Figure 3: Anti-symmetric kernel is obtained by negating and mirroring the trainable variables by the center point.
@@
Where $g_{s}$ represents the anti-symmetric continuous convolution kernel and $f$ represents the corresponding value at the position $x$, with the remaining symbols defined similarly to the definitions in Equation 1. The antisymmetric kernel $g_{s}$ satisfies the constraint condition $g_{s}(x)=-g_{s}(-x)$. For two distinct points $x,y\in P_{n}$, the interparticle force between these two points can be formulated as:
$F_{xy}$ $=(f_{x}+f_{y})(-g_{s}(x-y))$ (4)
$=-F_{yx}.$
Then Equation 4 can further lead to the validity of the following equation:
$\int_{x\in P_{n}}\int_{y\in P_{n}}F_{xy}=0.$ (5)
The antisymmetric design of ASCC ensures compliance with momentum conservation laws in data-driven learning, but this strong constraint also reduces its learning capacity, making the problem much harder to solve. In contrast, CConv has better fitting and learning capabilities but lacks constraints of physical laws. In the next section, we introduce our proposed Dual-pipeline architecture, which aims to build a neural network that integrates strong learning abilities while strictly adhering to physical principles.
@@
Figure 4: The architecture of the Attention-based Feature Fusion Module. Features from both pathways pass through two layers of CConv and are then concatenated. The concatenated features are subsequently input into a Particle Selector to determine the respective weights for fusion. The original features from both pathways are then combined with weighted fusion.
3.2 Dual-pipeline Architectrue
Our network architecture is illustrated in Figure 2. Instead of subjectively determining the relative significance of the two pathways, we employ an Attention-based Feature Fusion Module, allowing the network to autonomously learn the optimal connection between the two pathways, shown as Figure 4(a).
More precisely, our network comprises a main pathway and a constraint-guided pathway. Both pathways share a similar structure, each comprising a Type-aware Input Module and a sequence of continuous convolutions with an effective depth of five. The main pathway, which utilizes CConv as its kernel, is responsible for establishing a global context and stabilizing the fluid. The Constraint-guided pathway employs ASCC as its kernel, which enforces a strong constraint to ensure momentum conservation while learning fluid motion patterns. Before feeding the input into the convolutional sequence, we combine the result from the input module and a fully-connected layers by addition. In the second and fourth layers of the convolution sequence, we leverage the Feature Fusion Module to implement a residual connection concept [14]. We found that these designs improve accuracy and robustness.
Figure 5: The architecture of the Type-aware Input Module in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Main Path’s Type-aware Input Module replaces the ASCC kernel with CConv. The Particle Selector is the same as Figure 4(b).